Will a B(OH)₂ precipitate form if 50.0 mL of 0.03 M B₃X₂ is mixed with 350.0 mL of pH 11.5
Mg(OH)₂ solution at 25⁰C? Ksp of B(OH)₂ is 2.1 x 10⁻¹³.
1
Expert's answer
2020-12-30T04:02:37-0500

"B_3X_2 + 3Mg(OH)_2 \to 3B(OH)_2 + Mg_3X_2"


pH + pOH = 14

pOH = 14 - 11.5 = 2.5

pOH = -log(OH-) = 2.5

OH = 10-2.5

Therefore, [OH-] = 3.16 × 10-3M

1 molecule of "Mg(OH)_2" has 2 molecules of "OH^-". Therefore, the concentration of "Mg(OH)_2" is 6.32 × 10-3M



"B(OH)_{2(s)} \u2194 B^{2+} (aq) + 2OH^-_{(aq)}"; Ksp = 2.1 × 10-13


mol B2+ (aq) = (0.050 L sol)(0.03M) = 0.0015 mol B2+

[B2+] = 0.5 mol/0.400 L = 0.00375 M


mol OH- (aq) = 0.350 L sol)(0.00632M) = 0.00221 mol OH-

[Cl—] = 0.00221 mol/0.4000 L = 0.00553 M


Q = [B2+][OH-]2

 = (0.00375)(0.00553)2

 = 1.065 × 10-7

Q > Ksp

B(OH)2 will precipitate from solution.